OpenStudy (konradzuse):

Write and solve the differential equation that models the verbal statement. (Use k for the constant of proportionality. Use C for the constant of integration.) The rate of change of Q with respect to u is inversely proportional to the square of u. dQ/du = Q =

OpenStudy (chaise):

dq/du (is proportional to) 1/u^2 \[\frac{dQ}{dU} \alpha 1/u^2\] Thus: dq/du = k*(1/u^2) dq/du = k(1/u^2)

OpenStudy (konradzuse):

oic so we just integrated 1/u...

OpenStudy (konradzuse):

err nmvm I guess that's what we do now...

OpenStudy (anonymous):

it's separable... separate the variables...

OpenStudy (chaise):

Not quite, you don't integrate just yet. You make separate the variables, then it can be integrated. All the U's must be on one side of the equation, all the q's on the other.

OpenStudy (turingtest):

@chaise latex note: \propto\[\propto\]

OpenStudy (konradzuse):

dq = k(1/u^2)du?

OpenStudy (konradzuse):

q = k(-1/u) + C?

OpenStudy (konradzuse):

So is this correect? :)

OpenStudy (konradzuse):

yup guess so :P

OpenStudy (chaise):

I'm almost certain this is correct.

OpenStudy (konradzuse):

if my initial is 250g how would 250108 work? The practice version is like this too.... It says 147000 as the answer, but then says y=97....?

OpenStudy (chaise):

(ln(125/250))/1599=k Try using this for your k value, my bad. I think. :x

OpenStudy (konradzuse):

Because the initial quantity is 150 grams, y = 150ekt. Because the half-life is 1599 years, 10 = 20ek(1599) k = 1/1599 ln(1/2) . So, y = 150e[ln(1/2)/1599]t. When t = 1000, y = 150e[ln(1/2)/1599](1000) ≈ 97.24 g. When t = 3000, y ≈ 40.86 g.

OpenStudy (konradzuse):

this is 149976 lool

OpenStudy (konradzuse):

150e[ln(1/2)/1599](1000)

OpenStudy (konradzuse):

I don't get it..

OpenStudy (chaise):

http://www.wolframalpha.com/input/?i=250e%5E%28%28ln%281%2F2%29%2F1599%5D%281000%29%29 After 1000 years there will be 162 grams left. This makes sense, because after 1599 years there will be 125 grams left.

OpenStudy (konradzuse):

ah everything was e^ huh LOL :(

OpenStudy (konradzuse):

Stupid crap :(.